OpenStudy (anonymous):

write in the form y=a(x-h)^2+k y=-x^2+3x-2

OpenStudy (lgbasallote):

first group the x.. y = (-x^2 + 3x) - 2 now factor out -1 to make x positive y = -(x^2 - 3x) - 2 now complete the square

OpenStudy (anonymous):

is it y=-1(x-3/2)^2 +1/4

OpenStudy (anonymous):

need help :(

OpenStudy (lgbasallote):

i think you mean -3/2 @waterineyes :)

OpenStudy (anonymous):

Oh sorry..

OpenStudy (anonymous):

hehehe

OpenStudy (anonymous):

\[y=−x^22+3x−2=−(x^2−3x)−2=−((x)^2−3x+(\frac{3}{2})^2−(\frac{3}{2})^2)−2\] \[y=−((x−\frac{3}{2})^2−\frac{9}{4})−2=−(x−\frac{3}{2})^2+\frac{9}{4}−2=−(x−\frac{3}{2})^2+\frac{1}{4}\] \[y = −(x−\frac{3}{2})^2+\frac{1}{4}\]